OpenStudy (anonymous):

Calculate the ratio of effusion rates for methane (CH4) and nitrogen

OpenStudy (jfraser):

use graham's law of effusion \[\frac{R_2}{R_1} = \sqrt{\frac{MM_1}{MM_2}}\]